OpenStudy (anonymous):

solve the equation for exact solution over the interval of [0 - 2pi] : tan squared x + 3 = 0 .. help :)

OpenStudy (anonymous):

\[\tan ^{2}x+3=0\] is this the ?

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

it looks like there is no solution \[ \tan^{2} x=-3\] undefined

OpenStudy (anonymous):

really? .. thats my assignmnt hehehe..

OpenStudy (anonymous):

you can write tan^2 x = -3 tan x = (-3)^1\2 \[\tan x = \sqrt{3}i\]

OpenStudy (anonymous):

i think thats not it

OpenStudy (anonymous):

\[\tan x = \tan \frac{ \Pi }{ 3 } i\]

OpenStudy (anonymous):

\[x = \frac{ \Pi }{ 3 }i\]

OpenStudy (anonymous):

have you study about imaginary numbers

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

but we dont use imaginary in this part of the lesson i think

OpenStudy (anonymous):

we have to use imaginery number when there is -ve sign in the squareroot

OpenStudy (anonymous):

what do you say @camjan

OpenStudy (anonymous):

yes i know but there is really something wrong.. try another one ill just post another pls answer tnx